Grade 11Mechanics

An ideal gas with adiabatic constant ¥is heated at constant pressure. It absorbs q amount of heat. Determine the fraction of heat absorb in raising the internal energy and performing work.

Abhay
7 Years agoGrade 11

1 Answer

Ashutosh
7 Years ago
At constant pressure heat absorb by the system is q=ncpdT
 And internal energy change ∆u=ncvdT
Hence fraction ∆u/q=ncvdT/ncpdT=cv/co=1/¥
From first law of thermodynamics 
 
W=q-∆u
So fraction, w/q=q-∆u/q=1-∆u/q=1-1/¥.
It is a required answer.